An investment, which is worth 145,925 dollars and has an expected return of 3.6 percent, is expected to pay fixed annual cash flows for a given amount of time. The first annual cash flow is expected later today and the last annual cash flow is expected in 9 years from today. What is the present value of the annual cash flow that is expected in 2 years from today?
Answer #1

Annual cash flow=145925*3.6%/(1-1/1.036^10)*1/1.036=17021.982

Present value=17021.982/1.036^2=15859.54